OpenStudy (anonymous):

Show that the function sinx+cosx is of period 2pi. Also prove that sinx+cosx=sqrt(2)sin(x+pi/4)

OpenStudy (anonymous):

have no idea what to do with this one :-(

OpenStudy (anonymous):

\[sinx+cosx=\sqrt2\sin(x+\frac{ \pi }{ 4 })\]

OpenStudy (anonymous):

RHS: use \[\sin(A+B)=\sin(A)\cos(B)+\cos(A)\sin(B)\]

OpenStudy (anonymous):

what about that period Q

OpenStudy (anonymous):

well, \[\cos\theta=\cos(\theta+2\pi)\] so, start from let \[\cos\theta=\cos(\theta+A)\] solve for A

OpenStudy (anonymous):

would that make sense?

OpenStudy (anonymous):

sorry but I need to go somewhere i'll get back to u on this :)

OpenStudy (anonymous):

I am back!!

OpenStudy (anonymous):

sinx cos pi/4 + cosx sin pi/4 what's next?

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

help plz..... ??? where is everybody